OpenStudy (aravindg):

11. A cart with a mass M=1/2 kg is connected by a string to a weight of mass m= 200 g. At the initial moment the cart moves to the Left along a horizontal plane at a speed Vo =7 ms-1. Find the magnitude and Direction of the cart, the place where it Will be and the distance covered by It in t=5s.

OpenStudy (vishweshshrimali5):

from where u take all these questions

OpenStudy (anonymous):

first of all this is mechanics and not math but we'll call the initial speed of the cart Vo=7 m/s as there is no friction and the weight is not in the air the mass have no meaning for us if it was in the air we would have used newtons second low and say ma=mg - T (T is the tension in the string) and with the cart Ma=T and we would have found the acceleration of the system. as its not the case we will just use kinematics: so as I said Vo=7 (m/s) , a=0 (m/s^2) , Xo=0 (m) . from this data we get that the carts speed stay the same through out the whole ride --> V(5) = 7 (m/s). the distant is: X = Xo + Vo*t - 2at^2 --> put the data in and you'll get X = 7*5 = 35 (m) to left of where it started.